The probability density function of a random variable X is f (x) = C | x |; Find C,

and the value of xo such that ,

FX(xo)=3/4 where F is the CDF.


1
Expert's answer
2021-06-08T01:31:05-0400

Solution:

"f(x)=C|x|, \forall |x|\le a
\\f(x)=\{Cx, x\ge a\ge0
\\ \ \ \ \ \ \ \ \ \ \ \ \ \ -Cx,-a\le x<0"

Now, "\int_{-a}^0Cx \ dx+\int_0^{a}Cx \ dx=1"

"\Rightarrow C.\dfrac {x^2}2|_{-a}^0+C.\dfrac {x^2}2|_{0}^a=1"

"\Rightarrow C.\dfrac {a^2}2+C.\dfrac {a^2}2=1
\\ \Rightarrow C.a^2=1
\\ \Rightarrow C=\dfrac 1{a^2}"

We have to assume a value of "a" as it is not given, say, "a=1"

Then, "C=1"

So, "f(x)=|x|, \forall|x|\le 1"

Now, CDF"=F_X(x_0)=\int_{-1}^{x_0}(-x)dx+\int_{{x_0}}^{1}(x)dx=\frac34"

"\dfrac{-x^2}{2}|_{-1}^{x_0}+[\dfrac{x^2}{2}|_{x_0}^{1}]=\dfrac34"

"\dfrac{-{x_0}^2}{2}+\dfrac12+\dfrac12-\dfrac{{x_0}^2}{2}=\dfrac34"

"-{x_0}^2=\dfrac34-1
\\-{x_0}^2=-\dfrac14
\\x_0=\pm\dfrac12"